OpenStudy (anonymous):

The sides of a triangle are in the ratio of 1:2:2 and the perimeter of the triangle is 15 feet. The length of the shortest side is? 2.5 feet 3 feet 5 feet 7.5 feet

OpenStudy (campbell_st):

use the unitary method... which requires you to find the value of 1 unit in the ratio. add the parts in the ratio.....1 + 2 + 2 = the take the perimeter and divide but the number of parts in the ratio. this will be the value of 1 part ... and the length of the shorter side... hope it makes sense

OpenStudy (anonymous):

So 3 feet? 1+2+2=5/15=3?

OpenStudy (campbell_st):

that's correct